OpenStudy (anonymous):

To find the Calorie Density (D) in calories per ounce of a food that contains c calories and weighs w ounces is given by: D = C/W Eight ounces of fat-free milk contains 84 calories. Find the calorie density of the fat-free milk.

OpenStudy (anonymous):

C = 84, W = 8. Compute D = C/W = 84/8

OpenStudy (anonymous):

Thank you pgpilot326

OpenStudy (anonymous):

you're welcome

OpenStudy (kingofhou):

For the fat-free milk, c = 84 calories and w = 8 ounces. Plug in the values in the given equation to get: \[D = \frac{ 84 }{ 5 } = 10.5\] calories per ounce. The calorie density of the fat-free milk is 10.5 calories per ounce.